1. Sodium has mass number 22.9 and density 970 kg/m^3.

Find a) free electron concentration.

b) fermi wave number

C) fermi energy and velocity.


1
Expert's answer
2021-09-26T18:48:50-0400

a)

"n=\frac{\rho N_A}M=2.5\cdot 10^{28}~m^{-3},"

b)

"k=(3\pi^2 n)^{\frac 13}=9\cdot 10^9~m^{-1},"

c)

"E=\frac{h^2k^2}{8\pi^2m_e}=3.1~eV,"

"v=\frac{hk}{2\pi m_e}=10^6~\frac ms."